# Thermodynamic computing out of equilibrium

### v2

@@ III.2 Programming the computer
To program the computer we adjust the parameters ${\bm{\theta}}$ using a genetic algorithm instructed to minimize a loss function $\phi$. The loss is constructed from the $M$ samples (11) or (12) for all output neurons, and is evaluated for $K$ different sets of inputs. This evaluation requires $KM$ dynamical trajectories in reset-sampling mode, and $K$ dynamical trajectories in serial-sampling mode (loss functions are specified in Section III.3 and Section III.4). We consider a population of $P=50$ thermodynamic computers, each of which is initialized with random parameters $\theta_{i}\sim{\mathcal{N}}(0,10^{-2})$. We evaluate $\phi$ for each computer, and select the 5 computers associated with the smallest values of $\phi$. These 5 are cloned and mutated to produce a new population of 50 computers. Mutations are done by adding to each parameter of each computer a Gaussian random number $\epsilon\sim C^{-1/2}{\mathcal{N}}(0,10^{-2})$. The term $C$ is the fan-in. It is equal to 1 for biases $b_{i}$ or weights $f_{i}$; it is equal to $N_{j}$ for weights $W_{ij}$, where $N_{j}$ is the number of connections entering neuron $j$; and it is equal to $(N_{i}+N_{j})/2$ for weights $J_{ij}$. Scaling mutations by the fan-in ensures that changes to all neuron inputs are of similar scale even if the network is strongly heterogenous Whitelam et al. (2022) (similar scaling ideas were used when developing efficient implementations of gradient descent LeCun et al. (1996)). Values of $\phi$ for the new population are calculated, the best 5 are selected, and so on.
@@
We developed an efficient implementation of a genetic algorithm for GPU. In reset-sampling mode, each step of the genetic algorithm requires the evaluation of $PKM$ dynamical trajectories of length $t_{\rm f}$ ($KM$ trajectories per computer to construct the loss function $\phi$, with $P$ computers in the genetic population). In serial-sampling mode, each step of the genetic algorithm requires the evaluation of the $PK$ dynamical trajectories of length $t_{\rm f}=Mt_{\rm obs}$. All trajectories can be evaluated in parallel on a set of GPUs.
@@
III.3 Learning a nonlinear function
@@
$\phi\equiv K^{-1}\sum_{j=1}^{K}\left(y_{0}(z_{j})-y(z_{j})\right)^{2}.$ (14)
@@
is the output of the thermodynamic computer, given the input $z$, averaged over $M$ samples (either in reset-sampling mode (11) or serial-sampling mode (12)). To integrate Eq. (9) we use a first-order Euler scheme with timestep $\Delta t=10^{-3}$. We take $M=10^{3}$ for the purposes of training. The $f_{i}$ are parameters trained by the genetic algorithm.
@@
In Fig. 6(a) we show the loss as a function of evolutionary time for the two models, in reset-sampling mode. The linear model fails to train – it cannot express a nonlinear function of the input variable – while the nonlinear model learns steadily, reaching a small value of the loss. Panel (b) shows the output functions learned by the two models: the nonlinear model has learned a good approximation of the target cosine function. The intrinsic noise of the computer is visible in the output, but for $M=10^{3}$ samples, for each value of $z$, the mean output signal of the computer exceeds the scale of the noise by a considerable margin.
Panels (c) and (d) of Fig. 6 show the sampled neuron outputs as a function of the neuron inputs (the inputs being all signals in to the neuron, excepting the thermal noise) at the designated observation time, $t_{\rm f}=1$. Consistent with the considerations of Section II, the nonlinear model possesses a nonlinear finite-time activation function, indicating that we are beyond the nonlinear threshold observation time. The resulting nonlinear activation function explains the computer’s ability to learn an arbitrary nonlinear function. Panel (d) confirms that the quadratic-neuron computer is at finite times a linear model, as expected.
@@
Training is done using $M=10^{3}$ samples for each value of the input $z$, but the trained computer can be used with fewer than $M$ samples if desired. In Fig. 7 we show the output of the trained thermodynamic computer, as a function of the input $z$, for a range of values of $M$. The number of samples can be chosen in order to achieve a required precision.
@@
Training in reset-sampling mode results in a thermodynamic computer programmed to express the target function at a prescribed observation time $t_{\rm f}=1$. In Fig. 8 we show the output of the computer at a range of observation times. The output of the computer varies as a function of time, and is equal to the target function only at the prescribed observation time. The output of the computer in equilibrium (corresponding to the long-time limit) is considerably different to the target function. In this example, therefore, the programmed thermodynamic computer operates far from equilibrium.
@@
The model thermodynamic computer can also be trained in serial-sampling mode, as shown in Fig. 9. In this case the thermodynamic computer produces a trajectory whose stationary state approximates the target function over a range of observation times. The requirement of training is that, given an input $z$, the output of the computer, averaged over $M=10^{3}$ samples taken at evenly-spaced times, from $t_{\rm obs}=0.2$ to time $t_{\rm f}=Mt_{\rm obs}=200$, be as close as possible to the target function, $y_{0}(z)$, Eq. (13). In the top image of panel (c) we show the output of the trained computer, averaged over $M=10^{3}$ time slices, up to various $t_{\rm f}$. The output of the trained computer remains close to the target function for a wide range of observation times. The lower image in panel (c) shows the difference between the computer output and the target function: after an initial transient (which is included in the serial-sampling average (12)), the computer achieves a steady-state output that approximates the target function. This steady state may correspond to a true thermodynamic equilibrium, but we did not require this: the training requirement was only that the computer achieve a given output when averaged over a specific time interval.
@@
The color maps in Fig. 9(d) show similar quantities for the computer trained in reset-sampling mode (see Figs. 68). In this mode the computer was trained to approximate the target function using $M=10^{3}$ samples taken at a single observation time, $t_{\rm f}=1$ (see Eq. (11)). As shown by the color plots (derived from $10^{3}$ samples from the trained computer taken at observation time $t_{\rm f}$), the output of the computer approximates the target function only at time $t_{\rm f}=1$.
III.4 Classifying MNIST
@@
Figure 10: Training a simulation model of a thermodynamic computer to classify MNIST. The computer, which consists of a 3-layer network of quadratic-quartic $(1,0,1)$ neurons, is trained in reset-sampling mode, using $M=10^{3}$ samples taken at observation time $t_{\rm f}=1$. (a) Loss (18) as a function of evolutionary time $n$. (b) Training-set classification accuracy during training. (c) Test-set classification accuracy of the trained computer, as a function of the number of samples $M$ generated by the computer (each taken at observation time $t_{\rm f}=1$). (d) For a single digit, an 8, we show the probability distribution, taken over $10^{5}$ samples, of the computer’s per-sample class predictions, Eq. (19). The mean value (16) of each distribution, which is the value used for classification, is indicated at the top of the panel. The correct distribution is shown in green, the others in shades of blue. (e) As panel (d), but for a misclassified digit, a 5. (f) The class predictions (16) of the computer, upon being shown the indicated digit, for various observation time $t_{\rm f}$. The computer is trained to classify the digit at an observation time $t_{\rm f}=1$ (vertical dotted line).
@@
As in Section III.3, we take $\beta\equiv(k_{\rm B}T)^{-1}=10$, so that the neuron energy scale is 10 times that of the thermal energy, i.e. $J_{2,4}/k_{\rm B}T=10$. We consider a 3-layer thermodynamic computer with quadratic-quartic $(1,0,1)$ neurons. Each layer has 32 neurons. Each neuron in the input layer couples to all the pixels $I_{i}$ of an MNIST digit via Eq. (7). The output layer of 32 neurons is used to construct the computer’s prediction for the class of MNIST digit ${\bm{I}}_{j}$, via the 10 expressions
$y^{(C)}(\bm{I})\equiv\sum_{i\in{\rm outputs}}f_{i}^{(C)}\langle x_{i}({\bm{I}})\rangle_{\rm r}.$ (16)
Here $C\in[0,9]$ is the class index, and the $f_{i}^{(C)}$ are 320 parameters that will be trained by genetic algorithm. Recall that the reset-sampling average is specified by Eq. (11).
We train the computer in reset-sampling mode, with $M=10^{3}$ samples taken at observation time $t_{\rm f}=1$. For the loss function we choose the cross-entropy between the class probabilities predicted by the thermodynamic computer and the ground-truth labels. The probability $p^{(C)}$ that a given digit ${\bm{I}}_{j}$ is of class $C$ is obtained by applying a softmax transformation to (16),
$p^{(C)}({\bm{I}}_{j})=\frac{\exp\left[y^{(C)}({\bm{I}}_{j})\right]}{\sum_{C^{\prime}=0}^{9}\exp\left[y^{(C^{\prime})}({\bm{I}}_{j})\right]},$ (17)
@@
$\phi=-\frac{1}{K}\sum_{j=1}^{K}\sum_{C=0}^{9}\hat{p}^{(C)}_{k}({\bm{I}}_{j})\ln{p}^{(C)}_{k}({\bm{I}}_{j}).$ (18)
Here $\hat{p}$ is the ground-truth label, unity if ${\bm{I}}_{j}$ is of class $C$ and zero otherwise. The sum is taken over all $K=60$,000 training samples, i.e. we use full-batch learning.
This cross-entropy is used only in training: the expressions (17) and (18) are not intended to be implemented in hardware. Once trained, the computer can be used to do classification, which we do intend to be implementable using analog hardware. Classification is done by measuring which of the 10 quantities (16) is largest, when the computer is connected to a given digit. If the thermodynamic computer is realized in hardware, each of the outputs (16) could be connected to a tree of comparators and a multiplexer. This additional hardware could be used to determine the identity of the quantity with the largest value, and hence the computer’s predicted class for the digit.
@@
In Fig. 10(a) we show the loss (18) as a function of evolutionary time $n$ as the computer is trained to classify MNIST. The computer learns steadily under the action of the genetic algorithm. Panel (b) shows the corresponding training-set classification accuracy (which can be observed but is not used during training). Panel (c) shows the test-set classification accuracy of the trained computer, as a function of the number of samples $M$ (see Eq. (11)). Training was done using $M=10^{3}$ samples per digit, but the trained computer can be run with considerably fewer samples ($\approx 20$) without significant loss of accuracy. Thus while the training procedure involved the generation of more than $10^{13}$ simulated trajectories, the trained computer (which is designed to be implemented in hardware) can afford to generate as few as 20 trajectories in order to classify individual digits with reasonable accuracy.
@@
In Fig. 10(d) we show the output of the trained computer when presented with a single digit, an 8, which it correctly classifies. We plot the probability distribution, taken over $10^{5}$ samples, of the computer’s per-sample predictions
$Y^{(C)}(\bm{I})\equiv\sum_{i\in{\rm outputs}}f_{i}^{(C)}x_{i}({\bm{I}},t_{\rm f}).$ (19)
The mean value of each distribution, corresponding to Eq. (16), is indicated at the top of the panel. The correct distribution is shown in green, with the others in shades of blue. Panel (e) shows similar data for a digit, a 5, that is misclassified as a 6 by the computer.
Figure 11: Values of all neurons $x_{i}$, for a single trajectory of the 3-layer thermodynamic computer, when presented with an MNIST digit. The top panels are averaged over all neurons in the indicated layer. The computer’s outputs (the layer-3 neurons) are still evolving at the observation time $t_{\rm f}=1$, showing that the computer operates out of equilibrium.
Figure 12: MNIST test-set accuracy of the trained 3-layer thermodynamic computer at a range of energy scales (the parameters of the thermodynamic computer are held fixed, and the noise strength is varied). The computer is trained at the energy scale $J=10k_{\rm B}T$.
In Fig. 10(f) we plot the value of the trained computer’s 10 predictions (16) upon being shown the indicated digit. The computer is trained to classify the digit at an observation time $t_{\rm f}=1$. In this case the computer has attained a steady-state dynamics at the specified observation time, but this is not a general phenomenon: when presented with other digits, the computer’s neurons are still evolving at $t_{\rm f}=1$. In Fig. 11 we show the values of all neurons $x_{i}$ for a single trajectory of the computer, as a function of time, when presented with an MNIST digit different to the one used in Fig. 10(f). The top panels of Fig. 11 are averaged over all neurons in the indicated layer. The computer’s outputs are still evolving at $t_{\rm f}=1$, showing that the computer operates out of equilibrium.
III.5 Role of noise in the thermodynamic computer
@@
The ability of the trained computer to operate at different noise scales depends on the nature of the problem it is trained for and its architecture. In Fig. 12 we show the test-set accuracy of the trained thermodynamic computer at a range of energy scales (the parameters of the thermodynamic computer are held fixed, and the noise strength is varied). The computer, trained at the energy scale $J=10k_{\rm B}T$, performs essentially as well when subjected to noise comparable in scale to its own energy scale ($J=k_{\rm B}T$). This result shows the ability of the computer to operate reliably throughout the regime characteristic of thermodynamic computing, and shows its output to be robust to small changes in noise level (which might occur if a device becomes hot during computation). For sufficiently large noise levels ($J\lesssim k_{\rm B}T$) the computer’s performance begins to decline.
Figure 13: (a) Loss associated with a simulated thermodynamic computer trained at zero noise to express a cosine function and operated at various finite noise levels. (b) MNIST Test-set accuracy of a 3-layer thermodynamic computer trained at zero noise to classify MNIST and operated at various finite noise levels (blue). The orange line shows the same thing for a 9-layer computer. The vertical black dashed line denotes the energy scale $J=k_{\rm B}T$.
@@
In Fig. 13(a) we show the loss associated with a thermodynamic computer trained to express the cosine function at zero noise ($J/k_{\rm B}T\to\infty$) and tested at a wide range of energy scales. The output values of the computer change continuously with energy scale, and so while it can tolerate the very low noise levels characteristic of classical computing, it breaks down for the larger noise levels characteristic of thermodynamic computing. In this case, the computer must be trained in the presence of noise to operate in the presence of noise.
In Fig. 13(b) we show the MNIST test-set accuracy of a 3-layer thermodynamic computer, trained at zero noise to classify MNIST, when operated at various finite noise levels (blue). This case is different to that of panel (a); here, the accuracy changes very little as the noise is increased, even up to the noise levels characteristic of thermodynamic computing. The mean outputs of the computer change continuously with noise, and so therefore does the loss, but classification accuracy depends on the relative values of the computer’s outputs, not their absolute values, and so is less strongly affected. This is a convenient feature, because training is much less expensive in the deterministic zero-noise limit. There, we used backpropagation through time Werbos (1990) on a single trajectory of the computer (for each digit), and training takes about 30 minutes on a single GPU instead of 24 hours on 96 GPUs. For some applications, therefore, it is possible to train a thermodynamic computer at zero noise, and operate it at noise levels comparable to the energy scale of the device.
The orange line in Fig. 13(b) shows the same thing for a 9-layer thermodynamic computer trained at zero noise. This model is more accurate than the 3-layer one for small noise values, but fails more rapidly as noise is increased, and fails before we reach the large noise levels characteristic of thermodynamic computing. The reason for this failure is likely the diminishing signal-to-noise ratio, in the interior of the computer, as we move away from the input layer. This effect can be seen in the 3-layer computer shown in Fig. 11: the signal scale, and hence the signal-to-noise ratio, diminishes from input to output layers. The comparison of the the two lines in Fig. 13(b) indicates that robustness to noise levels different to those seen during training is an architecture-dependent phenomenon, and suggests that the layered architecture characteristic of classical neural networks may not be an optimal design for a nonlinear thermodynamic computer.
@@
IV Conclusions
Classical computing aims to overcome or suppress thermal fluctuations, while thermodynamic computing uses thermal fluctuations. Most thermodynamic computing algorithms proposed to date require the computer to reach thermodynamic equilibrium, and arrange for the Boltzmann distribution of the computer’s degrees of freedom to encode the outcome of a desired calculation. Although a powerful approach, one difficulty it encounters is that equilibration times for physical systems can span a broad range of timescales; a second is that not every calculation can be expressed in an obvious way by the Boltzmann distribution.
An alternative is to arrange for a thermodynamic computer to perform arbitrary nonlinear calculations at specified observation times, whether or not the computer has come to thermal equilibrium. In this paper we have presented the design for such a computer, a “thermoneural network”, and have programmed a simulation model of the computer by genetic algorithm to express a nonlinear function and to classify MNIST digits. The core of our design is a thermodynamic neuron, a fluctuating classical degree of freedom confined by a quartic potential. (In general, the confining potential need not be exactly quartic, but it must be higher-order than quadratic and thermodynamically stable; we can consider Eq. (1) to represent a Maclaurin expansion, in powers of $x$, of an arbitrary nonlinear thermodynamic circuit.) The neuron’s activity is a nonlinear function of its input, in and out of equilibrium, and so interacting graphs of such neurons can function as universal approximators, in and out of equilibrium. Thermodynamic neural networks can be programmed by genetic algorithm to perform nonlinear computations analogous to those performed by conventional neural networks, at specified times, regardless of whether the system has reached equilibrium. Classical computers operate on a clock, and we have shown that thermodynamic computers can do so, too.
The thermodynamic computer design presented here uses several elements that have already been deployed in hardware. For instance, RLC circuits can encode fluctuating degrees of freedom that interact via bilinear couplings Aifer et al. (2024); Melanson et al. (2023). However, the degrees of freedom of existing thermodynamic computers possess quadratic self-interactions, which result in linear activation functions and cannot be used to construct a universal approximator. Nonlinear inductors or capacitors might be used to induce higher-order self interactions within RLC circuits Kuh and Hajj (1971); Chua (1980). Alternatively, thermodynamic computers could be built from superconducting circuits, in which case Josephson junctions can provide nonlinear inductance Ray and Crutchfield (2023). If such designs can be implemented, this paper shows that the resulting nonlinear computer can be programmed, e.g. by genetic algorithm. The trained computer would evolve by thermal fluctuations to perform a specified computation at a specified observation time, whether or not the computer has attained thermodynamic equilibrium.
Acknowledgements.
This work was done at the Molecular Foundry, supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. C.C. was supported by a Francqui Fellowship of the Belgian American Educational Foundation, and by the US DOE Office of Science Scientific User Facilities AI/ML project “A digital twin for spatiotemporally resolved experiments”.

### v3

@@ Abstract
To program the computer we adjust the parameters $\bm{\theta}$} using a genetic algorithm instructed to minimize a loss function $\phi$. The loss is constructed from the $M$ samples (S11) or (S12) for all output neurons, and is evaluated for $K$ different sets of inputs. This evaluation requires $KM$ dynamical trajectories in reset-sampling mode, and $K$ dynamical trajectories in serial-sampling mode (loss functions are specified in Section S3.3 and Section S3.4). We consider a population of $P=50$ thermodynamic computers, each of which is initialized with random parameters $\theta_{i}\sim{\mathcal{N}}(0,10^{-2})$. We evaluate $\phi$ for each computer, and select the 5 computers associated with the smallest values of $\phi$. These 5 are cloned and mutated to produce a new population of 50 computers. Mutations are done by adding to each parameter of each computer a Gaussian random number $\epsilon\sim C^{-1/2}{\mathcal{N}}(0,10^{-2})$. The term $C$ is the fan-in. It is equal to 1 for biases $b_{i}$ or weights $f_{i}$; it is equal to $N_{j}$ for weights $W_{ij}$, where $N_{j}$ is the number of connections entering neuron $j$; and it is equal to $(N_{i}+N_{j})/2$ for weights $J_{ij}$. Scaling mutations by the fan-in ensures that changes to all neuron inputs are of similar scale even if the network is strongly heterogenous Whitelam et al. (2022) (similar scaling ideas were used when developing efficient implementations of gradient descent LeCun et al. (1996)). Values of $\phi$ for the new population are calculated, the best 5 are selected, and so on.
@@
We developed an efficient implementation of a genetic algorithm for GPU. In reset-sampling mode, each step of the genetic algorithm requires the evaluation of $PKM$ dynamical trajectories of length $t_{\rm f}$ (KMKM trajectories per computer to construct the loss function $\phi$, with $P$ computers in the genetic population). In serial-sampling mode, each step of the genetic algorithm requires the evaluation of the $PK$ dynamical trajectories of length $t_{\rm f}=Mt_{\rm obs}$. All trajectories can be evaluated in parallel on a set of GPUs.
@@
S3.3 Learning a nonlinear function
@@
$\phi\equiv K^{-1}\sum_{j=1}^{K}\left(y_{0}(z_{j})-y(z_{j})\right)^{2}.$ (S14)
@@
$y(z)\equiv\sum_{i\in{\rm outputs}}f_{i}\langle x_{i}(z)\rangle_{{\rm r},{\rm s}}$ (S15)
is the output of the thermodynamic computer, given the input $z$, averaged over $M$ samples (either in reset-sampling mode (S11) or serial-sampling mode (S12)). To integrate Eq. (S9) we use a first-order Euler scheme with timestep $\Delta t=10^{-3}$. We take $M=10^{3}$ for the purposes of training. The $f_{i}$ are parameters trained by the genetic algorithm.
@@
In Fig. S6(a) we show the loss as a function of evolutionary time for the two models, in reset-sampling mode. The linear model fails to train – it cannot express a nonlinear function of the input variable – while the nonlinear model learns steadily, reaching a small value of the loss. Panel (b) shows the output functions learned by the two models: the nonlinear model has learned a good approximation of the target cosine function. The intrinsic noise of the computer is visible in the output, but for $M=10^{3}$ samples, for each value of $z$, the mean output signal of the computer exceeds the scale of the noise by a considerable margin.
Panels (c) and (d) of Fig. S6 show the sampled neuron outputs as a function of the neuron inputs (the inputs being all signals in to the neuron, excepting the thermal noise) at the designated observation time, $t_{\rm f}=1$. Consistent with the considerations of Section S2, the nonlinear model possesses a nonlinear finite-time activation function, indicating that we are beyond the nonlinear threshold observation time. The resulting nonlinear activation function explains the computer’s ability to learn an arbitrary nonlinear function. Panel (d) confirms that the quadratic-neuron computer is at finite times a linear model, as expected.
@@
Training is done using $M=10^{3}$ samples for each value of the input $z$, but the trained computer can be used with fewer than $M$ samples if desired. In Fig. S7 we show the output of the trained thermodynamic computer, as a function of the input $z$, for a range of values of $M$. The number of samples can be chosen in order to achieve a required precision.
@@
Training in reset-sampling mode results in a thermodynamic computer programmed to express the target function at a prescribed observation time $t_{\rm f}=1$. In Fig. S8 we show the output of the computer at a range of observation times. The output of the computer varies as a function of time, and is equal to the target function only at the prescribed observation time. The output of the computer in equilibrium (corresponding to the long-time limit) is considerably different to the target function. In this example, therefore, the programmed thermodynamic computer operates far from equilibrium.
@@
The model thermodynamic computer can also be trained in serial-sampling mode, as shown in Fig. S9. In this case the thermodynamic computer produces a trajectory whose stationary state approximates the target function over a range of observation times. The requirement of training is that, given an input $z$, the output of the computer, averaged over $M=10^{3}$ samples taken at evenly-spaced times, from $t_{\rm obs}=0.2$ to time $t_{\rm f}=Mt_{\rm obs}=200$, be as close as possible to the target function, $y_{0}(z)$, Eq. (S13). In the top image of panel (c) we show the output of the trained computer, averaged over $M=10^{3}$ time slices, up to various $t_{\rm f}$. The output of the trained computer remains close to the target function for a wide range of observation times. The lower image in panel (c) shows the difference between the computer output and the target function: after an initial transient (which is included in the serial-sampling average (S12)), the computer achieves a steady-state output that approximates the target function. This steady state may correspond to a true thermodynamic equilibrium, but we did not require this: the training requirement was only that the computer achieve a given output when averaged over a specific time interval.
@@
The color maps in Fig. S9(d) show similar quantities for the computer trained in reset-sampling mode (see Figs. S6S8). In this mode the computer was trained to approximate the target function using $M=10^{3}$ samples taken at a single observation time, $t_{\rm f}=1$ (see Eq. (S11)). As shown by the color plots (derived from $10^{3}$ samples from the trained computer taken at observation time $t_{\rm f}$), the output of the computer approximates the target function only at time $t_{\rm f}=1$.
S3.4 Classifying MNIST
@@
Figure S10: Training a simulation model of a thermodynamic computer to classify MNIST. The computer, which consists of a 3-layer network of quadratic-quartic $(1,0,1)$ neurons, is trained in reset-sampling mode, using $M=10^{3}$ samples taken at observation time $t_{\rm f}=1$. (a) Loss (S18) as a function of evolutionary time $n$. (b) Training-set classification accuracy during training. (c) Test-set classification accuracy of the trained computer, as a function of the number of samples $M$ generated by the computer (each taken at observation time $t_{\rm f}=1$). (d) For a single digit, an 8, we show the probability distribution, taken over $10^{5}$ samples, of the computer’s per-sample class predictions, Eq. (S19). The mean value (S16) of each distribution, which is the value used for classification, is indicated at the top of the panel. The correct distribution is shown in green, the others in shades of blue. (e) As panel (d), but for a misclassified digit, a 5. (f) The class predictions (S16) of the computer, upon being shown the indicated digit, for various observation time $t_{\rm f}$. The computer is trained to classify the digit at an observation time $t_{\rm f}=1$ (vertical dotted line).
@@
As in Section S3.3, we take $\beta\equiv(k_{\rm B}T)^{-1}=10$, so that the neuron energy scale is 10 times that of the thermal energy, i.e. $J_{2,4}/k_{\rm B}T=10$. We consider a 3-layer thermodynamic computer with quadratic-quartic $(1,0,1)$ neurons. Each layer has 32 neurons. Each neuron in the input layer couples to all the pixels $I_{i}$ of an MNIST digit via Eq. (S7). The output layer of 32 neurons is used to construct the computer’s prediction for the class of MNIST digit ${\bm{I}}_{j}$, via the 10 averaged class scores
$y^{(C)}(\bm{I})\equiv\sum_{i\in{\rm outputs}}f_{i}^{(C)}\langle x_{i}({\bm{I}})\rangle_{\rm r}.$ (S16)
Here $C\in[0,9]$ is the class index, and the $f_{i}^{(C)}$ are 320 parameters that will be trained by genetic algorithm. Recall that the reset-sampling average is specified by Eq. (S11).
We train the computer in reset-sampling mode, with $M=10^{3}$ samples taken at observation time $t_{\rm f}=1$. For the loss function we choose the cross-entropy between the class probabilities predicted by the thermodynamic computer and the ground-truth labels. The probability $p^{(C)}$ that a given digit ${\bm{I}}_{j}$ is of class $C$ is obtained by applying a softmax transformation to (S16),
$p^{(C)}({\bm{I}}_{j})=\frac{\exp\left[y^{(C)}({\bm{I}}_{j})\right]}{\sum_{C^{\prime}=0}^{9}\exp\left[y^{(C^{\prime})}({\bm{I}}_{j})\right]},$ (S17)
@@
$\phi=-\frac{1}{K}\sum_{j=1}^{K}\sum_{C=0}^{9}\hat{p}^{(C)}_{k}({\bm{I}}_{j})\ln{p}^{(C)}_{k}({\bm{I}}_{j}).$ (S18)
Here $\hat{p}$ is the ground-truth label, unity if ${\bm{I}}_{j}$ is of class $C$ and zero otherwise. The sum is taken over all K=60K=60,000 training samples, i.e. we use full-batch learning.
Figure S11: Values of all neurons $x_{i}$, for a single trajectory of the 3-layer thermodynamic computer, when presented with an MNIST digit. The top panels are averaged over all neurons in the indicated layer. The computer’s outputs (the layer-3 neurons) are still evolving at the observation time $t_{\rm f}=1$, showing that the computer operates out of equilibrium.
This cross-entropy is used only in training: the expressions (S17) and (S18) are not intended to be implemented in hardware. Once trained, the computer can be used to do classification, which we do intend to be implementable using analog hardware. Classification is done by measuring which of the 10 quantities (S16) is largest, when the computer is connected to a given digit. If the thermodynamic computer is realized in hardware, each of the outputs (S16) could be connected to a tree of comparators and a multiplexer. This additional hardware could be used to determine the identity of the quantity with the largest value, and hence the computer’s predicted class for the digit.
@@
In Fig. S10(a) we show the loss (S18) as a function of evolutionary time $n$ as the computer is trained to classify MNIST. The computer learns steadily under the action of the genetic algorithm. Panel (b) shows the corresponding training-set classification accuracy (which can be observed but is not used during training). Panel (c) shows the test-set classification accuracy of the trained computer, as a function of the number of samples $M$ (see Eq. (S11)). Training was done using $M=10^{3}$ samples per digit, but the trained computer can be run with considerably fewer samples ($\approx 20$) without significant loss of accuracy. Thus while the training procedure involved the generation of more than $10^{13}$ simulated trajectories, the trained computer (which is designed to be implemented in hardware) can afford to generate as few as 20 trajectories in order to classify individual digits with reasonable accuracy.
@@
In Fig. S10(d) we show the output of the trained computer when presented with a single digit, an 8, which it correctly classifies. We plot the probability distribution, taken over $10^{5}$ samples, of the computer’s per-sample predictions
$Y^{(C)}(\bm{I})\equiv\sum_{i\in{\rm outputs}}f_{i}^{(C)}x_{i}({\bm{I}},t_{\rm f}).$ (S19)
The mean value of each distribution, corresponding to Eq. (S16), is indicated at the top of the panel. The correct distribution is shown in green, with the others in shades of blue. Panel (e) shows similar data for a digit, a 5, that is misclassified as a 6 by the computer.
Figure S12: MNIST test-set accuracy of the trained 3-layer thermodynamic computer at a range of energy scales (the parameters of the thermodynamic computer are held fixed, and the noise strength is varied). The computer is trained at the energy scale $J=10k_{\rm B}T$.
In Fig. S10(f) we plot the value of the trained computer’s 10 predictions (S16) upon being shown the indicated digit. The computer is trained to classify the digit at an observation time $t_{\rm f}=1$. In this case the computer has attained a steady-state dynamics at the specified observation time, but this is not a general phenomenon: when presented with other digits, the computer’s neurons are still evolving at $t_{\rm f}=1$. In Fig. S11 we show the values of all neurons $x_{i}$ for a single trajectory of the computer, as a function of time, when presented with an MNIST digit different to the one used in Fig. S10(f). The top panels of Fig. S11 are averaged over all neurons in the indicated layer. The computer’s outputs are still evolving at $t_{\rm f}=1$, showing that the computer operates out of equilibrium.
S3.5 Role of noise in the thermodynamic computer
@@
The ability of the trained computer to operate at different noise scales depends on the nature of the problem it is trained for and its architecture. In Fig. S12 we show the test-set accuracy of the trained thermodynamic computer at a range of energy scales (the parameters of the thermodynamic computer are held fixed, and the noise strength is varied). The computer, trained at the energy scale $J=10k_{\rm B}T$, performs essentially as well when subjected to noise comparable in scale to its own energy scale ($J=k_{\rm B}T$). This result shows the ability of the computer to operate reliably throughout the regime characteristic of thermodynamic computing, and shows its output to be robust to small changes in noise level (which might occur if a device becomes hot during computation). For sufficiently large noise levels ($J\lesssim k_{\rm B}T$) the computer’s performance begins to decline.
Figure S13: (a) Loss associated with a simulated thermodynamic computer trained at zero noise to express a cosine function and operated at various finite noise levels. (b) MNIST Test-set accuracy of a 3-layer thermodynamic computer trained at zero noise to classify MNIST and operated at various finite noise levels (blue). The orange line shows the same thing for a 9-layer computer. The vertical black dashed line denotes the energy scale $J=k_{\rm B}T$.
@@
In Fig. S13(a) we show the loss associated with a thermodynamic computer trained to express the cosine function at zero noise ($J/k_{\rm B}T\to\infty$) and tested at a wide range of energy scales. The output values of the computer change continuously with energy scale, and so while it can tolerate the very low noise levels characteristic of classical computing, it breaks down for the larger noise levels characteristic of thermodynamic computing. In this case, the computer must be trained in the presence of noise to operate in the presence of noise.
In Fig. S13(b) we show the MNIST test-set accuracy of a 3-layer thermodynamic computer, trained at zero noise to classify MNIST, when operated at various finite noise levels (blue). This case is different to that of panel (a); here, the accuracy changes very little as the noise is increased, even up to the noise levels characteristic of thermodynamic computing. The mean outputs of the computer change continuously with noise, and so therefore does the loss, but classification accuracy depends on the relative values of the computer’s outputs, not their absolute values, and so is less strongly affected. This is a convenient feature, because training is much less expensive in the deterministic zero-noise limit. There, we used backpropagation through time Werbos (1990) on a single trajectory of the computer (for each digit), and training takes about 30 minutes on a single GPU instead of 24 hours on 96 GPUs. For some applications, therefore, it is possible to train a thermodynamic computer at zero noise, and operate it at noise levels comparable to the energy scale of the device.
The orange line in Fig. S13(b) shows the same thing for a 9-layer thermodynamic computer trained at zero noise. This model is more accurate than the 3-layer one for small noise values, but fails more rapidly as noise is increased, and fails before we reach the large noise levels characteristic of thermodynamic computing. The reason for this failure is likely the diminishing signal-to-noise ratio, in the interior of the computer, as we move away from the input layer. This effect can be seen in the 3-layer computer shown in Fig. S11: the signal scale, and hence the signal-to-noise ratio, diminishes from input to output layers. The comparison of the the two lines in Fig. S13(b) indicates that robustness to noise levels different to those seen during training is an architecture-dependent phenomenon, and suggests that the layered architecture characteristic of classical neural networks may not be an optimal design for a nonlinear thermodynamic computer.